1. Joni Shamrock Inc. has the following amounts reported in its general ledger at the end of the current year.

Organization costs $23,600
Trademarks 15,300
Discount on bonds payable 36,600
Deposits with advertising agency for ads to promote goodwill of company 11,600
Excess of cost over fair value of net identifiable assets of acquired subsidiary 76,600
Cost of equipment acquired for research and development projects;
       the equipment has an alternative future use
86,600
Costs of developing a secret formula for a product that is expected
       to be marketed for at least 20 years
82,200


(a)On the basis of this information, compute the total amount to be reported by Shamrock for intangible assets on its balance sheet at year-end.

Total amount reported for intangible assets $enter the total dollar amount to be reported for intangible assets

2.

The project is completed in 2020, and a successful patent is obtained. The R&D costs to complete the project are $122,000. The administrative and legal expenses incurred in obtaining patent number 472-1001-84 in 2020 total $18,500. The patent has an expected useful life of 5 years. Record these costs in journal entry form. Also, record patent amortization (full year) in 2020. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

(To record research and development costs)

(To record legal and administrative costs)

(To record one year’s amortization expense)

In 2021, the company successfully defends the patent in extended litigation at a cost of $59,200, thereby extending the patent life to December 31, 2028. What is the proper way to account for this cost? Also, record patent amortization (full year) in 2021. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

(To record legal cost of successfully defending patent)

(To record one year’s amortization)
Answer #1

Answer:

a.)

Trademarks    $15,300
Excess of cost over fair value of net identifiable assets of acquired subsidiary (Goodwill) $76,600
Total Intangible Assets $91,900

2.)

Particulars Debit ($) Credit ($)
Research and Development Expense $122,000
          Cash $122,000
(To record Research and Development costs)
Patents $18,500
          Cash $18,500
(To record legal and administrative costs)
Amortization Expense $3,700 =18500/5
         Patents $3,700
(To record one year's amortization expense)
Particulars Debit ($) Credit ($)
patents $59,200
Cash $59,200
(To record legal cost successfully defends the patent)
Patents amortization expense $9,250
Patents $9,250

Working Notes:

patents = $18500/5 = 3700
Amortization Expense:
(18500 - 3700)/8 = $1,850
$59,200/8 = $7,400
Total = $9,250